OpenStudy (anonymous):

16 is what % of 64 please show all your work dont explain because i have 15 minuts to turn this in an it is a big part of my gradeplease do it and show your work i will give you a medal and I will fan you

OpenStudy (anonymous):

hold on..

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

"is" means = and "of" means * 16=? * 64

OpenStudy (anonymous):

hold on actually

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

16 is 25% of 64 so 16 as a percentage of 64 = 16/64 * 100 = 25 %

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

16 is 25% of 64 so 16 as a percentage of 64 = 16/64 * 100 = 25 %

OpenStudy (anonymous):

k bye

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

ur very welcome....I really hope it helped